Use the cut-out strips to determine 2÷3.
1
Expert's answer
2020-09-28T19:58:53-0400

Start by showing strip a with 3 pieces.

Then consider 1/3 of what is coloured in black.




"\dfrac{2}{3}=\dfrac{1}{3}+\dfrac{1}{3}=2×\dfrac{1}{3}"


Show 2 cut-out strips of "\dfrac{1}{3}"



Big black cut-out strips determine "\dfrac{2}{3}" .